OpenStudy (anonymous):

Help, please? In(3x - 2) = In40

OpenStudy (anonymous):

X = 14

OpenStudy (anonymous):

Could you explain to me how you got that answer, please?

OpenStudy (anonymous):

Okay, you have 3x - 2 = 40 Isolate terms with x on the left hand side so add 2 to both sides. 3x + (2 - 2) = 40 + 2 so now you have: 3x = 42 divide both sides by 3 and you have x = 14